OpenStudy (anonymous):

Where does the graph of this function cross the x-axis? f(x)=x^2+14x+24/3x^2-11

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The function is this right? \[\Large f(x) = \frac{x^2+14x+24}{3x^2-11}\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

First replace f(x) with 0. Then multiply both sides by \(\Large 3x^2-11\) \[\Large f(x) = \frac{x^2+14x+24}{3x^2-11}\] \[\Large 0 = \frac{x^2+14x+24}{3x^2-11}\] \[\Large 0*(3x^2-11) = \frac{x^2+14x+24}{3x^2-11}*(3x^2-11)\] \[\Large 0 = x^2+14x+24\] \[\Large x^2+14x+24=0\] Do you see how to solve from here?

OpenStudy (anonymous):

Got it! (x+2)(x+12) right?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so if (x+2)(x+12) = 0 then either x+2 = 0 or x+12 = 0 solve each equation for x to get your 2 answers

OpenStudy (anonymous):

x=-2 and x=-12?

jimthompson5910 (jim_thompson5910):

both are correct

OpenStudy (anonymous):

All right, thank you so much!!

jimthompson5910 (jim_thompson5910):

you're welcome